I need to find the unknown of 2/3/10=1.8/r
I think it is r=27 or no solution, please help.

Found 2 solutions by richwmiller, checkley77:
Answer by richwmiller(17219)

Why would there be no solution/
And why would would it be 27?

Answer by checkley77(12844)

2/3/10=1.8/r
(2/3)*1/10=1.8/r
2/30=1.8/r
2r=1.8*30
2r=54
r=54/2
r=27 answer.
Proof:
(2/3)*1/10=1.8/27
2/30=1.8/27
54=54